OpenStudy (anonymous):

I have a problem with the question Evaluate sin 20π/ 6

OpenStudy (anonymous):

because 20*pi/6 = 20*180/6 =600 you have to subtract 600-360=240 that is sin(20*pi/6)=sin(240)

OpenStudy (anonymous):

wait, there is more

OpenStudy (anonymous):

Sin(240)=-sqrt(3)/2

OpenStudy (anonymous):

that's it

OpenStudy (anonymous):

I got it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Please if you dont... just ask

OpenStudy (anonymous):

but why do you subtration 360

OpenStudy (anonymous):

In one circle, you always sustract 360° in order for the angle to be from 0 to 360°. Sorry my english is bad and I cannot explain it better

OpenStudy (anonymous):

I can understand your anwser

OpenStudy (anonymous):

answer

OpenStudy (anonymous):

OK :)